OpenStudy (anonymous):

5 cm in 1 minute = how many meters in 1 week

OpenStudy (unanimoose):

Let's convert the minute into a week. 1 * 60 = 60 60 * 24 = 1440 1440 * 7 = 10080 So: 5cm * 10080 = 50,400

OpenStudy (unanimoose):

50,400 cm = 504 meters

OpenStudy (unanimoose):

Whoever runs 5cm in 1 minute should really see a specialist.